OpenStudy (yrelhan4):

number of ways in which score of 11 can be made from a throw of 3 persons each throwing a single die once?

OpenStudy (yrelhan4):

ok so what i did is x1 + x2 +x3 = 11, number of solutions=13c2 then we have to subtract, 3*(4c2) --> refer http://openstudy.com/users/yrelhan4#/updates/51320a88e4b0f0611bc045e6 but i am not getting the answer. where am i going wrong?

OpenStudy (yrelhan4):

correction: 3*(5c2)

OpenStudy (yrelhan4):

@shubhamsrg

OpenStudy (unklerhaukus):

641, 632, 551, 542, 533, 443,

OpenStudy (agent0smith):

Kinda tedious but you could do it by hand... to get 11 (hopefully i didn't miss any): 6, 4, 1 6, 3, 2 5, 5, 1 5, 4, 2 5, 3, 3 4, 4, 3

OpenStudy (yrelhan4):

are you sure you have to do it by hand? answer is 27. :O

OpenStudy (agent0smith):

Then you might have to account for the fact that each of those can happen in 3P3 ways, if it matters who throws what number.

OpenStudy (agent0smith):

Then you'll have to subtract a couple, since 5,3,3 is the same as 5,3,3

OpenStudy (unklerhaukus):

the sets with three different numbers has 6 possible orders, the sets with two different numbers have 3 possible orders, the sets with one repeated number have only one order . 6+6+3+6+3+3=

OpenStudy (yrelhan4):

hmmm. thank you. :)

OpenStudy (agent0smith):

It seems difficult to do with just straight formulas, given that your x1 x2 x3 values are limited between 1 and 6.

OpenStudy (yrelhan4):

i know. but we did a problem yesterday with those formulas. i think i am doing something wrong. i'll just give it a try and post the work if i get it.

OpenStudy (agent0smith):

6, 4, 1 .... 3P3 6, 3, 2 .... 3P3 5, 5, 1 .... (3P3)/2 5, 4, 2 .... 3P3 5, 3, 3 .... (3P3)/2 4, 4, 3 .... (3P3)/2 so 3*(3P3) + 3*(3P3)/2

OpenStudy (yrelhan4):

^ hmm. i'll note down that method too. :)

OpenStudy (yrelhan4):

ok i got what i was looking for. first since x,y,z>=1, the equation will be x+y+z=7.. number of solutions = n-1Cr-1=9C2 now referring to the method in the link i posted before (i dont know how to explain it), we have to subtract 3*(4C2). hence the answer.

OpenStudy (yrelhan4):

thank you everybody, i appreciate. :)

OpenStudy (shubhamsrg):

for that method, yep, we'll first account for 6>=x>=1 , 6>=y>=1 and 6>=z>=1, or 5>=x>=0 , 5>=y>=0 ,5>=z>=0 => (5-x) + (5-y) + (5-z) = 7 => C(9,2) = 36 from here, we will subtract cases when our X>=6, Y>=6 and Z>=6 => 3* C(3,2) = 9 -> ans= 36-9 = 27

OpenStudy (shubhamsrg):

make that 5>=X>=0 , 5>=Y>=0 ,5>=Z>=0 where X=x-1 Y=y-1 and Z=z-1

OpenStudy (yrelhan4):

yup.. i was a little wrong there.